One leg of a right triangle is twice the length of the other. The hypotenuse is 8 m long. Find the length of each leg. Report your answer to the nearest tenth of a meter.
Answer by checkley77(12844)

A^2+B^2=C^2
A^2+(2A)2=8^2
A^2+4A^2=64
5A^2=64
5A^2=64
A^2=64/5
A^2=12.8
A=SQRT12.8
A=3.6 METERS IS THE SHORT SIDE.
2*3.6=7.2 METERS IS THE LENGTH OF THE LONGER SIDE.
PROOF:
3.6^2+7.2^2=64
12.96+51.84=64
64.8~64